### v5

# Cohomology and deformations of left-symmetric Rinehart algebras

Abstract
We introduce a notion of left-symmetric Rinehart algebras, which is a generalization of the notion of left-symmetric algebras. The left multiplication gives rise to a representation of the corresponding sub-adjacent Lie-Rinehart algebra. We construct left-symmetric Rinehart algebras from $\mathcal{O}$-operators on Lie-Rinehart algebras. We extensively investigate representations of left-symmetric Rinehart algebras. Moreover, we construct a graded Lie algebra on the space of multi-derivations whose Maurer–Cartan elements characterize left-symmetric Rinehart algebras and study deformations of left-symmetric Rinehart algebras, which are controlled by the second cohomology class in the deformation cohomology. We also give the relationships between $\mathcal{O}$-operators and Nijenhuis operators on left-symmetric Rinehart algebras.
keywords:
left-symmetric Rinehart algebra, representation, graded Lie algebra, Maurer-Cartan element, cohomology, deformation, Nijenhuis operator.
[…]
1 Introduction
Left-symmetric algebras are algebras for which the associator
$(x,y,z):=(x\cdot y)\cdot z-x\cdot(y\cdot z)$
satisfies the identity $(x,y,z)=(y,x,z)$. These algebras appeared as early as 1896 in the work of Cayley [8] as rooted tree algebras. In the 1960s, they also arose from the study of several topics in geometry and algebra, such as convex homogenous cones [30], affine manifolds and affine structures on Lie groups [19, 27] and deformations of associative algebras [14]. In 2006, Burde [6] wrote an interesting survey showing the importance of left-symmetric algebras in many areas, such as vector fields, rooted tree algebras, vertex algebras, operad theory, deformation complexes of algebras, convex homogeneous cones, affine manifolds and left-invariant affine structures on Lie groups [6].
Left symmetric algebras are the underlying algebraic structures of non-abelian phase spaces of Lie algebras [21, 1], leading to a bialgebra theory of left-symmetric algebras [3]. They can also be seen as the algebraic structures behind the classical Yang-Baxter equations. Precisely, they provide a construction of solutions of the classical Yang-Baxter equations in certain semidirect product Lie algebra structures (that is, over the double spaces) induced by left-symmetric algebras [22, 2].
The notion of Lie-Rinehart algebras was introduced by J. Herz in [15] and further developed in [28, 29]. A a notion of (Poincaré) duality for this class of algebras was introduced in [16, 17]. Lie-Rinehart structures have been the subject of extensive studies, in relations to symplectic geometry, Poisson structures, Lie groupoids and algebroids and other kinds of quantizations (see [18, 20, 25, 24, 23, 26]). For further details and a history of the notion of Lie-Rinehart algebra, we refer the reader to [18]. Lie-Rinehart algebras have been investigated furthermore in [4, 7, 12, 11].
A left-symmetric algebroid is a geometric generalization of a left-symmetric algebra. See [25, 24, 23] for more details and applications. The notion of a Nijenhuis operator on a left-symmetric algebroid was introduced in [24], which could generate a trivial deformation. More details on deformations of left-symmetric algebras can be found in [31].
In this paper, we introduce a notion of left-symmetric Rinehart algebras, which is a generalization of a left-symmetric algebra and an algebraic version of left symmetric algebroids. The following diagram shows how left-symmetric Rinehart algebras fit in relation to Lie algebras, left-symmetric algebras and Lie-Rinehart algebras.
Lie algebrageneralizationLie-Rinehart algebraLeft-symmetric algebrageneralizationLeft-symmetric Rinehart
The paper is organized as follows. In Section 2, we recall some definitions concerning left-symmetric algebras and Lie-Rinehart algebra. In Section 3, we introduce the notion of left-symmetric Rinehart algebra and give some of its properties. As in the case of a left-symmetric algebras, one can obtain the sub-adjacent Lie-Rinehart algebra from a left-symmetric Rinehart algebra by using the commutator. The left multiplication gives rise to a representation of the sub-adjacent Lie-Rinehart algebra. We construct left-symmetric Rinehart algebras using $\mathcal{O}$-operators. Section 4 is devoted to the study of representations and cohomology of left-symmetric Rinehart algebra. In Section 5, we construct a graded Lie algebra whose Maurer-Cartan elements are left-symmetric Rinehart algebras which give rise to a coboundary operator. Section 6 is devoted to introduce the deformation cohomology associated to a left-symmetric Rinehart algebra, which controls the deformations. In Section 7, we introduce the notion of a Nijenhuis operator, which could generate a trivial deformation. In addition, we investigate some connection between $\mathcal{O}$-operators and Nijenhuis operators.
Throughout this paper all vector spaces are over a field $\mathbb{K}$ of characteristic zero.
2 Preliminaries
In this section, we briefly recall some basics of left-symmetric algebras and Lie-Rinehart algebras [6].
Definition 2.1.
A left-symmetric algebra is a vector space $L$ endowed with a linear map $\cdot:L\otimes L\longrightarrow L$ such that for any $x,y,z\in L$,
$(x,y,z)=(y,x,z),\;\;{\rm or}\;\;{\rm equivalently,}\;\;(x\cdot y)\cdot z-x\cdot(y\cdot z)=(y\cdot x)\cdot z-y\cdot(x\cdot z),$
where the associator $(x,y,z):=(x\cdot y)\cdot z-x\cdot(y\cdot z)$.
Let $\mathrm{ad}^{L}$ (resp. $\mathrm{ad}^{R}$) be the left multiplication operator (resp. right multiplication operator) on $L$ that is, i.e. $\mathrm{ad}^{L}(x)y=x\cdot y$ (resp. $\mathrm{ad}^{R}(x)y=y\cdot x$), for any $x,y\in L$. The following lemma is given in [6].
Lemma 2.2.
Let $(L,\cdot)$ be a left-symmetric algebra. The commutator $[x,y]=x\cdot y-y\cdot x$ defines a Lie algebra $L$, which is called the sub-adjacent Lie algebra of $L$. The algebra $L$ is also called a compatible left-symmetric algebra on the Lie algebra $L$. Furthermore, the map $ad^{L}:L\rightarrow\mathfrak{gl}(L)$ with $x\mapsto L_{x}$ gives a representation of the Lie algebra $(L,[\cdot,\cdot])$.
Definition 2.3.
Let $(L,\cdot)$ be a left-symmetric algebra and $M$ a vector space. A representation of $L$ on $M$ consists of a pair $(\rho,\mu)$, where $\rho:L\longrightarrow\mathfrak{gl}(M)$ is a representation of the sub-adjacent Lie algebra $L$ on $M$ and $\mu:L\longrightarrow\mathfrak{gl}(M)$ is a linear map satisfying:
$\rho(x)\circ\mu(y)-\mu(y)\circ\rho(x)=\mu(x\cdot y)-\mu(y)\circ\mu(x),\quad\forall~{}x,y\in L.$ (1)
The map $\rho$ is called a left representation and $\mu$ is a right representation. Usually, we denote a representation by $(M;\rho,\mu)$. Then $(L;\mathrm{ad}^{L},\mathrm{ad}^{R})$ is a representation of $(L,\cdot)$ which is called adjoint representation.
The cohomology complex for a left-symmetric algebra $(L,\cdot)$ with a representation $(M;\rho,\mu)$ is given as follows. The set of $(n+1)$-cochains is given by
$C^{n+1}(L,M)=\mathrm{Hom}(\wedge^{n}L\otimes L,M),\quad\forall n\geq 0.$ (2)
For all $\omega\in C^{n}(L,M)$, the coboundary operator $\delta:C^{n}(L,M)\longrightarrow C^{n+1}(L,M)$ is given by
$\delta\omega(x_{1},x_{2},\dots,x_{n+1})$ $=$ $\sum_{i=1}^{n}(-1)^{i+1}\rho(x_{i})\omega(x_{1},\dots,\widehat{x_{i}},\dots,x_{n+1})$
$\quad+\sum_{i=1}^{n}(-1)^{i+1}\mu(x_{n+1})\omega(x_{1},\dots,\widehat{x_{i}},\dots,x_{n},x_{i})$
$\quad-\sum_{i=1}^{n}(-1)^{i+1}\omega(x_{1},\dots,\widehat{x_{i}},\dots,x_{n},x_{i}\cdot x_{n+1})$
$\quad+\sum_{1\leq i<j\leq n}(-1)^{i+j}\omega([x_{i},x_{j}],x_{1},\dots,\widehat{x_{i}},\dots,\widehat{x_{j}},\dots,x_{n+1}).$
We then have the following lemma whose proof comes from a direct computation using identity (1).
Lemma 2.4 (See [5]).
The map $\delta$ satisfies $\delta^{2}=0$.
Definition 2.5.
A Lie-Rinehart algebra $L$ over an associative commutative algebra $A$ is a Lie algebra over $\mathbb{K}$ with an $A$-module structure and a linear map $\rho:L\rightarrow Der(A)$, such that the following conditions hold:
  1. 1.
    For all $a\in A$ and $x,y\in L$
    $\rho([x,y])=\rho(x)\rho(y)-\rho(x)\rho(y)\quad\text{and}\quad\rho(ax)=a\rho(x).$
  2. 2.
    The compatibility condition:
    $[x,ay]=\rho(x)ay+a[x,y],~{}~{}~{}\forall a\in A,x,y\in L.$ (3)
Let $(L,A,[\cdot,\cdot]_{L},\rho)$ and $(L^{\prime},A^{\prime},[\cdot,\cdot]_{L^{\prime}},\rho^{\prime})$ be two Lie-Rinehart algebras, then a Lie-Rinehart algebra homomorphism is defined as a pair of maps $(g,f)$, where the maps $f:L\rightarrow L^{\prime}$ and $g:A\rightarrow A^{\prime}$ are two algebra homomorphisms such that:
  • (1)
    $f(ax)=g(a)f(x)$ for all $x\in L$ and $a\in A,$
  • (2)
    $g(\rho(x)a)=\rho^{\prime}(f(x))g(a)$ for all $x\in L$ and $a\in A.$
Now, we recall the definition of module over a Lie-Rinehart algebra (for more details see [10]).
Definition 2.6.
Let $M$ be an $A$-module. Then $M$ is a module over a Lie-Rinehart algebra $(L,A,[\cdot,\cdot],\rho)$ if there exists a map $\theta:L\otimes M\rightarrow M$ such that:
  1. 1.
    $\theta$ is a representation of the Lie algebra $(L,[\cdot,\cdot])$ on $M$.
  2. 2.
    $\theta(ax,m)=a\theta(x,m)$ for all $a\in A,x\in L,m\in M.$
  3. 3.
    $\theta(x,am)=a\theta(x,m)+\rho(x)am$ for all $x\in L,a\in A,m\in M.$
We have the following lemma giving a characterization of of the $\theta$ which are representations.
Lemma 2.7.
The map $\theta$ is representation if and only if $L\oplus M$ is Lie-Rinehart algebra over $A$, where $[\cdot,\cdot]_{L\oplus M}$ and $\theta_{L\oplus M}$ are given by
$[x_{1}+m_{1},x_{2}+m_{2}]_{L\oplus M}$ $=$ $[x_{1},x_{2}]+\rho(x_{1})m_{2}-\rho(x_{2})m_{1},$
$\theta_{L\oplus M}(x_{1}+m_{1})$ $=$ $\theta(x_{1})$
for all $x_{1},x_{2}\in L$ and $m_{1},m_{2}\in M$.
3 Some basic properties of a left-symmetric Rinehart algebras
In this section, we introduce a notion of left-symmetric Rinehart algebras illustrated by some examples. As in the case of a left-symmetric algebra, we obtain the sub-adjacent Lie-Rinehart algebra from a left-symmetric Rinehart algebra using the commutator. In addition, we construct left-symmetric Rinehart algebras using $\mathcal{O}$-operators.
Definition 3.1.
A left-symmetric Rinehart algebra is a quadruple $(L,A,\cdot,\ell)$ where $(L,\cdot)$ is a left-symmetric algebra, $A$ is an associative commutative algebra and $\ell:L\rightarrow Der(A)$ a linear map such that the following conditions hold:
  1. 1.
    $L$ is an $A$-module.
  2. 2.
    For all $a\in A$ and $x,y\in L$
    $\ell(x\cdot y-y\cdot x)=\ell(x)\ell(y)-\ell(x)\ell(y)\quad\text{and}\quad\ell(ax)=a\ell(x).$
  3. 3.
    The compatibility conditions: for all $a\in A$ and $x,y\in L$
    $x\cdot(ay)=$ $\ell(x)ay+a(x\cdot y),$ (4)
    $(ax)\cdot y=$ $a(x\cdot y).$ (5)
Example 3.2.
It is clear that any left-symmetric algebra is a left-symmetric Rinehart algebra.
Example 3.3.
A Novikov Poisson algebra is a left-symmetric Rinehart algebra (see [32]).
Example 3.4.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and let $L\oplus A$be the direct sum of $L$ and $A$. Then $(L\oplus A,A,\cdot_{L\oplus A},\ell_{L\oplus A})$ is a left-symmetric Rinehart algebra, where the $\cdot_{L\oplus A}$ is defined by the following expression, for all $x_{1},x_{2}\in L$, $a_{1},a_{2}\in A$;
$(x_{1}+a_{1})\cdot_{L\oplus A}(x_{2}+a_{2})=x_{1}\cdot x_{2}+\ell(x_{1})(a_{2});$
and $\ell_{L\oplus A}:L\oplus A\rightarrow Der(A)$ is defined by $\ell_{L\oplus A}(a_{1}+x_{1})=\ell(x_{1})$. Indeed, it obvious that $(L\oplus A,\cdot_{L\oplus A})$ is a left-symmetric algebra, $\ell_{L\oplus A}$ is a representation of left-symmetric algebra $L\oplus A$ and $\ell_{L\oplus A}\in Der(A)$.
By direct calculation, we have $\ell_{L\oplus A}(b(x_{1}+a_{1}))=b\ell_{L\oplus A}(x_{1}+a_{1})$ for all $b,a_{1}\in A$ and $x_{1}\in L$. On the other hand, letting $x_{1},x_{2}\in L$ and $b,a_{1},a_{2}\in A$, we have
$(x_{1}+a_{1})\cdot_{L\oplus A}b(x_{2}+a_{2})=$ $(a_{1}+x_{1})\cdot_{L\oplus A}(bx_{2}+ba_{2})$
$=$ $x_{1}\cdot(bx_{2})+\ell(x_{1})(ba_{2})$
$=$ $b(x_{1}\cdot x_{2})+\ell(x_{1})b(x_{2})+\ell(x_{1})(b)a_{2}+b\ell(x_{1})(a_{2})$
$=$ $b\big{(}x_{1}\cdot x_{2}+\ell(x_{1})(a_{2})\big{)}+\ell(x_{1})b(x_{2})+\ell(x_{1})(b)a_{2}$
$=$ $b\big{(}(x_{1}+a_{1})\cdot_{L\oplus A}(x_{2}+a_{2})\big{)}+\ell_{L\oplus A}(x_{1}+a_{1})b(x_{2}+a_{2}).$
Moreover,
$b(x_{1}+a_{1})\cdot_{L\oplus A}(x_{2}+a_{2})=$ $(bx_{1}+ba_{1})\cdot_{L\oplus A}(x_{2}+a_{2})$
$=$ $(bx_{1})\cdot x_{2}+\ell(bx_{1})(a_{2})$
$=$ $b(x_{1}\cdot x_{2})+b\ell(x_{1})(a_{2})$
$=$ $b\big{(}x_{1}\cdot x_{2}+\ell(x_{1})a_{2}\big{)}$
$=$ $b\big{(}(x_{1}+a_{1})\cdot_{L\oplus A}(x_{2}+a_{2})\big{)}.$
Now we have the following theorem.
Theorem 3.5.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra. Then, $(L,A,[\cdot,\cdot],\ell)$ is a Lie-Rinehart algebra, denoted by $L^{C}$, called the sub-adjacent Lie-Rinehart algebra of $(L,A,\cdot,\ell)$.
Proof 3.6.
Since $(L,\cdot)$ is a left-symmetric algebra, we have that $(L,[\cdot,\cdot])$ is a Lie algebra. For any $a\in A$, by direct computations, we have
$[x,ay]$ $=$ $x\cdot(ay)-(ay)\cdot x=a(x\cdot y)+\ell(x)ay-a(y\cdot x)$
$=$ $a[x,y]+\ell(x)(a)y,$
which implies that $(L,A,[\cdot,\cdot],\ell)$ is a Lie-Rinehart algebra.
To see that the linear map $\ell:L\longrightarrow Der(A)$ is a representation, we only need to show that $\ell_{[x,y]}=[\ell_{x},\ell_{y}]_{Der(A)}$, which follows directly from the fact that $(L,\cdot)$ is a left-symmetric algebra. This ends the proof.
Definition 3.7.
Let $(L_{1},A_{1},\cdot_{1},\ell_{1})$ and $(L_{2},A_{2},\cdot_{2},\ell_{2})$ be two left-symmetric Rinehart algebras. A homomorphism of left-symmetric Rinehart algebras is a pair of two algebra homomorphisms $(f,g)$ where $f:L_{1}\longrightarrow L_{2}$ and $g:A_{1}\longrightarrow A_{2}$ such that:
$f(ax)=g(a)f(x),\>~{}g(\ell_{1}(x)a)=\ell_{2}(f(x))g(a),\quad\forall x,y\in L_{1},a\in A_{1}.$
The following proposition is immediate.
Proposition 3.8.
Let $(f,g)$ be a homomorphism of left-symmetric Rinehart algebras from $(L_{1},A_{1},\cdot_{1},\ell_{1})$ to $(L_{2},A_{2},\cdot_{2},\ell_{2})$. Then $(f,g)$ is also a Lie-Rinehart algebra homomorphism of the corresponding sub-adjacent Lie-Rinehart algebras.
Now we give the definition of an $\mathcal{O}$-operator.
Definition 3.9.
Let $(L,A,[\cdot,\cdot],\rho)$ be a Lie-Rinehart algebra and $\theta:L\longrightarrow End(M)$ be a representation over $M$. A linear map $T:M\longrightarrow L$ is called an $\mathcal{O}$-operator if for all $u,v\in M$ and $a\in A$ we have
$T(au)=aT(u),$ (6)
$[T(u),T(v)]=T\big{(}\theta(T(u))(v)-\theta(T(v))(u)\big{)}.$ (7)
Remark 3.10.
Consider the semidirect product Lie-Rinehart algebra
$(L\ltimes_{\theta}M,A,[\cdot,\cdot]_{L\ltimes_{\theta}M},\rho_{L\ltimes_{\theta}M}),$
where $\rho_{L\ltimes_{\theta}M}(x+u):=\rho(x)(u)$ and the bracket $[\cdot,\cdot]_{L\ltimes_{\theta}M}$ is given by
$[x+u,y+v]_{L\ltimes_{\theta}M}=[x,y]+\theta(x)(v)-\theta(y)(u).$
Any $\mathcal{O}$-operator $T:M\longrightarrow L$ gives a Nijenhuis operator $\tilde{T}=\left(\begin{array}[]{cc}0&T\\ 0&0\end{array}\right)$ on the Lie-Rinehart algebra $L\ltimes_{\theta}M$. More precisely, we have
$[\tilde{T}(x+u),\tilde{T}(y+v)]_{L\ltimes_{\theta}M}=\tilde{T}\Big{(}[\tilde{T}(x+u),y+v]_{L\ltimes_{\theta}M}+[x+u,\tilde{T}(y+v)]_{L\ltimes_{\theta}M}-\tilde{T}[x+u,y+v]_{L\ltimes_{\theta}M}\Big{)}.$
Fore more details on Nijenhuis operators and their applications the reader should consult [13].
Let $T:M\longrightarrow L$ be an $\mathcal{O}$-operator. Define the multiplication $\cdot_{T}$ on $M$ by
$u\cdot_{T}v=\theta(T(u))(v),\>\forall u,v\in M.$
We then have the following proposition.
Proposition 3.11.
With the above notations, $(M,A,\cdot_{T},\ell_{T}=\ell\circ T)$ is a left-symmetric Rinehart algebra, and the map $T$ is Lie-Rinehart algebra homomorphism from $(M,[\cdot,\cdot])$ to $(L,[\cdot,\cdot])$.
Proof 3.12.
It is easy to see that $(M,\cdot_{T})$ is a left-symmetric algebra. For any $a\in A$, using Definition 3.1 and equation (6) we have
$\ell_{M}(au)=\ell(T(au))=a\ell(T(u))=a\ell_{M}(u),$
Similarly, using Definition 2.6 we obtain
$(au)\cdot_{T}v$ $=$ $\theta(T(au))(v)=\theta aT(u))(v)=a\theta(T(u))(v),$
$u\cdot_{T}(av)$ $=$ $\theta(T(u))(av)=a\theta(T(u))(v)+\ell\circ T(u)(a)v.$
Thus, $(M,A,\cdot_{T},\ell_{M})$ is a left-symmetric Rinehart algebra. Let [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] be the sub-adjacent Lie bracket on $M$. Then we have
$T[u,v]=T(u\cdot_{T}v-v\cdot_{T}u)=T(\theta(T(u))(v)-\theta(T(v))(u))=[T(u),T(v)].$
So $T$ is a homomorphism of Lie algebras.
4 Representations of left-symmetric Rinehart algebras
In this section, we develop the notion of representations of a left-symmetric Rinehart algebra and give a cohomology theory with coefficients in a representation.
Definition 4.1.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $M$ be an $A$-module. A representation of $A$ on $M$ consists of a pair $(\rho,\mu)$, where $\rho$ is a representation of the sub-adjacent Lie-Rinehart algebra $(L,A,[\cdot,\cdot]^{C},\ell)$ and $\mu:L\rightarrow End(M)$ is a linear map, such that for all $x,y\in L$ and $\ m\in M$, we have
$\mu(ax)m=a\mu(x)m=\mu(x)(am)$
$\rho(x)\mu(y)-\mu(y)\rho(x)=\mu(x\cdot y)-\mu(y)\mu(x).$ (8)
We will denote this representation by $(M;\rho,\mu)$.
For a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$ and a representation $(M;\rho,\mu)$, the following proposition gives a construction of a left-symmetric Rinehart algebra called semidirect product and denoted by $L\ltimes_{\rho,\mu}M$.
Proposition 4.2.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $(M;\rho,\mu)$ a representation. Then, $(L\oplus M,A,\cdot_{L\oplus M},\ell_{L\oplus M})$ is a left-symmetric Rinehart algebra, where $\cdot_{L\oplus M}$ and $\ell_{L\oplus M}$ are given by
$(x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2})$ $=$ $x_{1}\cdot x_{2}+\rho(x_{1})m_{2}+\mu(x_{2})m_{1},$ (9)
$\ell_{L\oplus M}(x_{1}+m_{1})$ $=$ $\ell(x_{1}),$ (10)
for all $x_{1},x_{2}\in L$ and $\ m_{1},m_{2}\in M$.
Proof 4.3.
Let $(M;\rho,\mu)$ be a representation. It is straightforward to see that $(L\oplus M,A,\cdot_{L\oplus M})$ is a left-symmetric algebra. For any $x_{1},x_{2}\in L$ and $m_{1},m_{2}\in M$, we have
$(x_{1}+m_{1})\cdot_{L\oplus M}(a(x_{2}+m_{2}))$ $=$ $x_{1}\cdot(ax_{2})+\rho(x_{1})am_{2}+\mu(ax_{2})m_{1}$
$=$ $a(x_{1}\cdot x_{2})+\ell(x_{1})(ax_{2})+a\rho(x_{1})m_{2}$
$\qquad\qquad\>+\ell(x_{1})(am_{2})+a\mu(x_{2})m_{1}$
$=$ $a((x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2}))+\ell_{L\oplus M}(x_{1})(a)(x_{2}+m_{2}).$
On the other hand, we have
$(a(x_{1}+m_{1}))\cdot_{L\oplus M}(x_{2}+m_{2})$ $=$ $(ax_{1})\cdot x_{2}+\rho(ax_{1})m_{2}+\mu(x_{2})(am_{1})$
$=$ $a((x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2})).$
Therefore, $(L\oplus M,A,\cdot_{L\oplus M},\ell_{L\oplus M})$ is a left-symmetric Rinehart algebra.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $(M;\rho,\mu)$ be a representation.
Let $\rho^{*}:L\otimes M^{*}\rightarrow M^{*}$ and $\mu^{*}:M^{*}\otimes L\longrightarrow M^{*}$ be defined by
$\langle\rho^{*}(x)\xi,m\rangle=\ell(x)\langle\xi,m\rangle-\langle\rho(x)m,\xi\rangle\ \ \textrm{and}\ \ \langle\mu^{*}(x)\xi,m\rangle=-\langle\xi,\mu(x)m\rangle,$
where $M^{*}=Hom_{A}(M,A)$. Then, we have the following proposition.
Proposition 4.4.
With the above notations, we obtain that
  • $\rm(i)$
    $(M,\rho-\mu)$ is a representation of the sub-adjacent Lie-Rinehart algebra $(L,A,[\cdot,\cdot],\ell)$.
  • $\rm(ii)$
    $(M^{*},\rho^{*}-\mu^{*},-\mu^{*})$ is a representation be a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$.
Proof 4.5.
Since $(M;\rho,\mu)$ is a representation of the left-symmetric algebra $(L,A,\cdot,\ell)$, using Proposition 4.2 we have that $(L\oplus M,A,\cdot_{L\oplus M},\ell_{L\oplus M})$ is a left-symmetric Rinehart algebra. Consider its sub-adjacent Lie-Rinehart algebra $(L\oplus M,A,\cdot_{L\oplus M},[\cdot,\cdot]_{L\oplus M},\ell_{L\oplus M})$. We have
$[(x_{1}+m_{1}),(x_{2}+m_{2})]_{L\oplus M}$ $=$ $(x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2})-(x_{2}+m_{2})\cdot_{L\oplus M}(x_{1}+m_{1})$
$=$ $x_{1}\cdot x_{2}+\rho(x_{1})m_{2}+\mu(x_{2})m_{1}$
$-x_{2}\cdot x_{1}-\rho(x_{2})m_{1}-\mu(x_{1})m_{2}$
$=$ $[x_{1},x_{2}]^{C}+(\rho-\mu)(x_{1})(m_{2})+(\rho-\mu)(x_{2})(m_{1}).$
From Lemma 2.7 we deduce that $(M,\rho-\mu)$ is a representation of Lie-Rinehart algebra $L$ on $M$. This finishes the proof of (i).
For item (ii), it is clear that $\rho^{*}-\mu^{*}$ is just the dual representation of the representation $\theta=\rho-\mu$ of the sub-adjacent Lie-Rinehart algebra of $L$. We can directly check that $-\mu^{*}(ax)\xi=-a\mu^{*}(x)\xi=-\mu^{*}(x)(a\xi)$. For any $x,y\in L$, $\xi\in M^{*}$ and $m\in M$ we have
$-\langle(\rho^{*}-\mu^{*})(x)\mu^{*}(y)\xi,m\rangle+\langle\mu^{*}(y)(\rho^{*}-\mu^{*})(x)\xi,m\rangle$
$=-\langle\rho^{*}(x)\mu^{*}(y)\xi,m\rangle+\langle\mu^{*}(x)\mu^{*}(y)\xi,m\rangle+\langle\mu^{*}(y)\rho^{*}(x)\xi,m\rangle-\langle\mu^{*}(y)\mu^{*}(x)\xi,m\rangle$
$=\ell(x)\langle\xi,\mu(y)m\rangle-\langle\xi,\mu(y)\rho(x)m\rangle+\langle\xi,\mu(y)\mu(x)m\rangle$
$\qquad-\ell(x)\langle\xi,\mu(y)m\rangle+\langle\xi,\rho(x)\mu(y)m\rangle-\langle\xi,\mu(x)\mu(y)m\rangle$
$=\langle\xi,\mu(x\cdot y)m\rangle-\langle\xi,\mu(y)\mu(x)m\rangle+\langle\xi,\mu(y)\mu(x)m\rangle-\langle\xi,\mu(x)\mu(y)m\rangle$
$=\langle(-\mu^{*}(x\cdot y)-\mu^{*}(y)\mu^{*}(x))\xi,m\rangle.$
Therefore $(M^{*},\rho^{*}-\mu^{*},-\mu^{*})$ is a representation of $L$.
Corollary 4.6.
With the above notations, we have
  • $\rm(i)$
    The left-symmetric Rinehart algebras $L\ltimes_{\rho,\mu}M$ and $L\ltimes_{\rho-\mu,0}M$ have the same sub-adjacent Lie-Rinehart algebra $L\ltimes_{\rho-\mu}M.$
  • $\rm(ii)$
    The left-symmetric Rinehart algebras $L\ltimes_{\rho^{*},0}M^{*}$ and $L\ltimes_{\rho^{*}-\mu^{*},-\mu^{*}}M^{*}$ have the same sub-adjacent Lie-Rinehart algebra $L\ltimes_{\rho^{*}}M^{*}$.
Let $(M;\rho,\mu)$ be a representation of a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$. In general, $(M^{*},\rho^{*},\mu^{*})$ is not a representation. But we have the following proposition.
Proposition 4.7.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $(M;\rho,\mu)$ be a representation. Then the following conditions are equivalent:
  • $\rm(1)$
    $(M;\rho-\mu,-\mu)$ is a representation of $(L,A,\cdot,\ell)$.
  • $\rm(2)$
    $(M^{*},\rho^{*},\mu^{*})$ is a representation of $(L,A,\cdot,\ell)$.
  • $\rm(3)$
    $\mu(x)\mu(y)=\mu(y)\mu(x)$ for all $x,y\in L.$
5 The Matsushima-Nijenhuis bracket for left-symmetric Rinehart algebras
In this section we construct a graded Lie algebra whose Maurer-Cartan elements are left-symmetric Rinehart algebras which give rise to a coboundary operator.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebras. A multiderivation of degree $n$ is a multilinear map $P\in\mathrm{Hom}(\Lambda^{n}L\otimes L,L)$ such that, for every $a\in A$, $x_{i}\in L$ and $i\in\{1,2,\dotsc,n+1\}$, we have
$P(x_{1},\dots,ax_{i},\dots,x_{n},x_{n+1})=aP(x_{1},\dots,x_{i},\dots,x_{n},x_{n+1}),$ (11)
$P(x_{1},\cdots,x_{n},ax_{n+1})=aP(x_{1},\cdots,x_{n},x_{n+1})+\Xi_{P}(x_{1},\cdots,x_{n})(a)x_{n+1},$ (12)
where $\Xi_{P}:L^{\otimes n}\to Der(A)$ is called the symbol map. The space of all multiderivations of degree $n$ will be denoted by $\mathfrak{D}^{n}(L)$. Set $\mathfrak{D}^{*}(L)=\displaystyle\oplus_{n\geq-1}\mathfrak{D}^{n}(L)$ with $\mathfrak{D}^{-1}(L)=L$, the space of multiderivations on $L$.
A permutation $\sigma\in\mathbb{S}_{n}$ is called an $(i,n-i)$-unshuffle if $\sigma(1)<\cdots<\sigma(i)$ and $\sigma(i+~{}1)<~{}\cdots<\sigma(n)$. If $i=0$ and $i=n$, we assume $\sigma=Id$. The set of all $(i,n-i)$-unshuffles will be denoted by $\mathbb{S}_{(i,n-i)}$. The notion of an $(i_{1},\cdots,i_{k})$-unshuffle and the set $\mathbb{S}_{(i_{1},\cdots,i_{k})}$ are defined similarly.
Let $P\in\mathfrak{D}^{m}(L)$ and $Q\in\mathfrak{D}^{n}(L)$. We define the Matsushima–Nijenhuis bracket $[\cdot,\cdot]_{MN}:~{}\mathfrak{D}^{m}(L)\times\mathfrak{D}^{n}(L)\to\mathfrak{D}^{m+n}(L)$ by
$[P,Q]_{MN}=P\diamond Q-(-1)^{mn}Q\diamond P,$
where
$P\diamond Q(x_{1},x_{2},\cdots,x_{m+n+1})$
$=\sum_{\sigma\in\mathbb{S}_{(m,1,n-1)}}(-1)^{\sigma}P(Q(x_{\sigma(1)},\cdots,x_{\sigma(m+1)}),x_{\sigma(m+2)},\cdots,x_{\sigma(m+n)},x_{m+n+1})$
$\quad+(-1)^{mn}\sum_{\sigma\in\mathbb{S}_{(n,m)}}(-1)^{\sigma}P(x_{\sigma(1)},\cdots,x_{\sigma(n)},Q(x_{\sigma(n+1)},x_{\sigma(n+2)},\cdots,x_{\sigma(m+n)},x_{m+n+1})).$
Theorem 5.1.
With the above notations, we have
  • (i)
    The pair $(\mathfrak{D}^{*}(L),[\cdot,\cdot]_{MN})$ is a graded Lie algebra.
  • (ii)
    There is a one-to-one correspondence between the set of Maurer-Cartan elements of the graded Lie algebra $(\mathfrak{D}^{*}(L),[\cdot,\cdot]_{MN})$ and left-symmetric Rinehart algebra structures on $L$.
Proof 5.2.
(i) We begin by check that the Matsushima-Nijenhuis bracket is well defined. For $P\in\mathfrak{D}^{m}(L)$ and $Q\in\mathfrak{D}^{n}(L)$, by a direct calculation, we have
$[P,Q]_{MN}(ax_{1},x_{2},\cdots,x_{m+n+1})$
$=aP\diamond Q(x_{1},x_{2},\cdots,x_{m+n+1})-(-1)^{mn}aQ\diamond P(x_{1},x_{2},\cdots,x_{m+n+1})$
$+\sum_{\sigma\in\mathbb{S}_{(m-1,1,n-1)}}(-1)^{\sigma}\Xi_{Q}(x_{\sigma(2)},\cdots,x_{\sigma(m+1)})(a)P(x_{1},x_{\sigma(m+2)},\cdots,x_{\sigma(m+n)},x_{m+n+1})$
$+(-1)^{mn}\sum_{\sigma\in\mathbb{S}_{(n-1,1,m-1)}}(-1)^{\sigma}\Xi_{P}(x_{\sigma(2)},\cdots,x_{\sigma(n+1)})(a)Q(x_{1},x_{\sigma(n+2)},\cdots,x_{\sigma(m+n)},x_{m+n+1})$
$-(-1)^{mn}\sum_{\sigma\in\mathbb{S}_{(n-1,1,m-1)}}(-1)^{\sigma}\Xi_{P}(x_{\sigma(2)},\cdots,x_{\sigma(n+1)})(a)Q(x_{1},x_{\sigma(n+2)},\cdots,x_{\sigma(m+n)},x_{m+n+1})$
$-\sum_{\sigma\in\mathbb{S}_{(m-1,1,n-1)}}(-1)^{\sigma}\Xi_{Q}(x_{\sigma(2)},\cdots,x_{\sigma(m+1)})(a)P(x_{1},x_{\sigma(m+2)},\cdots,x_{\sigma(m+n)},x_{m+n+1})$
$=a[P,Q]_{MN}(x_{1},x_{2},\cdots,x_{m+n+1}),$
which implies that
$[P,Q]_{MN}(ax_{1},x_{2},\cdots,x_{m+n+1})=a[P,Q]_{MN}(x_{1},x_{2},\cdots,x_{m+n+1}).$
It is straightforward to check that $[P,Q]_{MN}$ is skew-symmetric with respect to its first $m+n$ arguments. Thus $[P,Q]_{MN}$ is $A$-linear with respect to its first $m+n$ arguments.
On the other hand, following a straightforward calculation, we have
$[P,Q]_{MN}(x_{1},x_{2},\cdots,ax_{m+n+1})$ $=$ $a[P,Q]_{MN}(x_{1},x_{2},\cdots,x_{m+n+1})$
$+\Xi_{[P,Q]_{MN}}(x_{1},x_{2},\cdots,x_{m+n})(a)x_{m+n+1},$
where the symbol map $\Xi_{[P,Q]_{MN}}$ is given by
$\Xi_{[P,Q]_{MN}}(x_{1},x_{2},\cdots,x_{m+n})(a)$
$=$ $\sum_{\sigma\in\mathbb{S}_{(m,1,n-1)}}(-1)^{\sigma}\Xi_{P}(Q(x_{\sigma(1)},\cdots,x_{\sigma(m+1)}),x_{\sigma(m+2)},\cdots,x_{\sigma(m+n)}))(a)$
$+$ $\sum_{\sigma\in\mathbb{S}_{(n,1,m-1)}}(-1)^{\sigma}\Xi_{Q}(P(x_{\sigma(1)},\cdots,x_{\sigma(n+1)}),x_{\sigma(n+2)},\cdots,x_{\sigma(m+n)})(a)$
$+$ $(-1)^{mn}\sum_{\sigma\in\mathbb{S}_{(m,n)}}(-1)^{\sigma}\Xi_{P}(x_{\sigma(1)},\cdots,x_{\sigma(n)})(\Xi_{Q}(x_{\sigma(n+1)},\cdots,x_{\sigma(n+m)}))(a)$
$+$ $\sum_{\sigma\in\mathbb{S}_{(m,n)}}(-1)^{\sigma}\Xi_{Q}(x_{\sigma(1)},\cdots,x_{\sigma(m)})(\Xi_{P}(x_{\sigma(m+1)},\cdots,x_{\sigma(m+n)}))(a).$
Thus $[P,Q]_{MN}\in\mathfrak{D}^{m+n}(L)$.
It was shown in [9] that the Matsushima-Nijenhuis bracket provides a graded Lie algebra structure on the graded vector space $\oplus_{n\geq 1}Hom(\Lambda^{n-1}L\otimes L,L)$. Therefore, $(\mathfrak{D}^{*}(L),[\cdot,\cdot]_{MN})$ is a graded Lie algebra.
(ii) Let $\pi\in\mathfrak{D}^{1}(L)$, we have
$\pi(ax_{1},x_{2})=a\pi(x_{1},x_{2}),\quad\pi(x_{1},ax_{2})=a\pi(x_{1},x_{2})+\Xi_{\pi}(x_{1})(a)x_{2},\quad\forall~{}x_{1},x_{2}\in L.$
In addition, we can easily check that
$[\pi,\pi]_{MN}(x_{1},x_{2},x_{3})$ $=$ $2(\pi(\pi(x_{1},x_{2}),x_{3})-\pi(\pi(x_{2},x_{1}),x_{3})$
$-\pi(x_{1},\pi(x_{2},x_{3}))+\pi(x_{2},(x_{1},x_{3}))).$
Thus $(L,A,\pi,\Xi_{\pi})$ is a left-symmetric Rinehart algebra if and only if $[\pi,\pi]_{MN}=0$.
Remark 5.3.
The cohomology of left-symmetric algebras first appeared in the unpublished paper of Y. Matsushima. Then A. Nijenhuis constructed a graded Lie bracket, which produces the cohomology theory for left-symmetric algebras. Thus the aforementioned graded Lie bracket is usually called the Matsushima–Nijenhuis bracket.
Let $(L,A,\pi,\ell)$ be a left-symmetric Rinehart algebra. According to Theorem 5.1, we have $[\pi,\pi]_{MN}=0$. Using the graded Jacobi identity, we get a coboundary operator $\delta:\mathfrak{D}^{n-1}(L)\to\mathfrak{D}^{n}(L)$, by putting
$\delta(P)=(-1)^{n-1}[\pi,P]_{MN},\quad\forall P\in\mathfrak{D}^{n-1}(L).$ (13)
By straightforward computation, we obtain
Proposition 5.4.
For any $P\in\mathfrak{D}^{n-1}(L)$, we have
$\delta P(x_{1},x_{2},\cdots,x_{n+1})$ $=$ $\sum_{i=1}^{n}(-1)^{i+1}\pi(x_{i},P(x_{1},x_{2},\cdots,\hat{x_{i}},\cdots,x_{n+1}))$
$+\sum_{i=1}^{n}(-1)^{i+1}\pi(P(x_{1},x_{2},\cdots,\hat{x_{i}},\cdots,x_{n},x_{i}),x_{n+1})$
$-\sum_{i=1}^{n}(-1)^{i+1}P(x_{1},x_{2},\cdots,\hat{x_{i}},\cdots,x_{n},\pi(x_{i},x_{n+1}))$
$+\sum_{1\leq i<j\leq{n}}(-1)^{i+j}P(\pi(x_{i},x_{j})-\pi(x_{j},x_{i}),x_{1},\cdots,\hat{x_{i}},\cdots,\hat{x_{j}},\cdots,x_{n+1})$
for all $\ x_{i}\in L,i=1,2\cdots,n+1$ and $\Xi_{\delta P}$ is given by
$\Xi_{P}(x_{1},x_{2},\cdots,x_{n})$ $=$ $\sum_{i=1}^{n}(-1)^{i+1}[\Xi_{\pi}(x_{i}),\Xi_{P}(x_{1},x_{2},\cdots,\hat{x_{i}},\cdots,x_{n})]$ (15)
$+\sum_{1\leq i<j\leq n}(-1)^{i+j}\Xi_{P}(\pi(x_{i},x_{j})-\pi(x_{j},x_{i}),x_{1},\cdots,\hat{x_{i}},\cdots,\hat{x_{j}},\cdots,x_{n})$
$+\sum_{i=1}^{n}(-1)^{i+1}\Xi_{\pi}(P(x_{1},x_{2},\cdots,\hat{x_{i}},\cdots,x_{n},x_{i})).$
Definition 5.5.
The cochain complex $(\mathfrak{D}^{*}(L)=\oplus_{n\geq 0}\mathfrak{D}^{n}(L),\delta)$ is called the deformation complex of the left-symmetric Rinehart algebra $L$. The corresponding $k$-th cohomology group, which we denote by $H^{k}(L)$, is called the $k$-th deformation cohomology group.
6 Deformation of left-symmetric Rinehart algebra
We investigate in this section a deformation theory of left-symmetric Rinehart algebras. But first let us introduce some notation. For a left-symmetric Rinehart algebra $(L,A,\mathfrak{m},\ell)$ we will denote the left-symmetric multiplication ``"``"`$\cdot$"` ` ⋅ " by $\mathfrak{m}$ in the sequel of the paper. Let $\mathbb{K}[[t]]$ be the formal power series ring in one variable $t$ and coefficients in $\mathbb{K}$. Let $L[[t]]$ be the set of formal power series whose coefficients are elements of $L$ (note that $L[[t]]$ is obtained by extending the coefficients domain of $\mathbb{K}[[t]]$ from $\mathbb{K}$ to $L$). Thus, $L[[t]]$ is a $\mathbb{K}[[t]]$-module.
6.1 Formal deformations
Definition 6.1.
A deformation of a left-symmetric Rinehart algebra $(L,A,\mathfrak{m},\ell)$ is a $\mathbb{K}[[t]]$-bilinear map
$\mathfrak{m}_{t}:L[[t]]\otimes L[[t]]\rightarrow L[[t]]$
which is given by $\mathfrak{m}_{t}(x,y)=\displaystyle\sum_{i\geq 0}t^{i}\mathfrak{m}_{i}(x,y),$ where $\mathfrak{m}_{0}=\mathfrak{m}$ and the $\mathfrak{m}_{i}\in\mathfrak{D}^{1}(L)$ satisfy the condition $[\mathfrak{m}_{t},\mathfrak{m}_{t}]_{MN}=0$.
Note that $\mathfrak{m}_{t}$ is a $1$-degree multiderivation of $A$ with symbol $\Xi_{\mathfrak{m}_{t}}:L\to Der(A)$ given by
$\Xi_{\mathfrak{m}_{t}}=\sum_{i\geq 0}t^{i}\Xi_{\mathfrak{m}_{i}}.$
Moreover, since $[\mathfrak{m}_{t},\mathfrak{m}_{t}]_{MN}=0$, it corresponds to a left symmetric Lie Rinehart algebra structure. In particular, it yields a $t$-parameterized family of products $\mathfrak{m}_{t}:L\otimes L\to L$ and a family of maps $\ell_{t}:L\rightarrow Der(A),$ which satisfy the following identities for all $x,y\in L$:
$\mathfrak{m}_{t}(x,y)=$ $x\cdot y+\displaystyle\sum_{i\geq 1}t^{i}\mathfrak{m}_{i}(x,y),$
$\ell_{t}(x)=$ $\ell(x)+\displaystyle\sum_{i\geq 1}t^{i}\Xi_{\mathfrak{m}_{i}}(x).$
The $t$-parametrized family $(L,A,\mathfrak{m}_{t},\ell_{t})$ is called a $1$-parameter formal deformation of $(L,A,\mathfrak{m},\ell)$ generated by $\mathfrak{m}_{1},\cdots,\mathfrak{m}_{m}\in\mathfrak{D}^{1}(L)$.
Let $(L,A,\mathfrak{m}_{t},\ell_{t})$ be a deformation of $\mathfrak{m}$. Then, for all $a\in A,x,y,z\in L$
$\mathfrak{m}_{t}(\mathfrak{m}_{t}(x,y),z)-\mathfrak{m}_{t}(x,\mathfrak{m}_{t}(y,z))=\mathfrak{m}_{t}(\mathfrak{m}_{t}(y,x),z)-\mathfrak{m}_{t}(y,\mathfrak{m}_{t}(x,z)).$ (16)
$\mathfrak{m}_{t}(ax,y)=a\mathfrak{m}_{t}(x,y),$ (17)
$\mathfrak{m}_{t}(x,ay)=a\mathfrak{m}_{t}(x,y)+\ell_{t}(x)ay$ (18)
The identities (17)–(18), mean that $\mathfrak{m}_{i}\in\mathfrak{D}^{1}(L)$. Comparing the coefficients of $t^{n}$ for $n\geq 0$ in equation (16), we get the following:
$\sum_{i+j=n}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z))=0.$ (19)
For $n=1$, equation (19) implies
$\mathfrak{m}_{1}(\mathfrak{m}(x,y),z)+\mathfrak{m}(\mathfrak{m}_{1}(x,y),z)-\mathfrak{m}_{1}(x,\mathfrak{m}(y,z))-\mathfrak{m}(x,\mathfrak{m}_{1}(y,z))$
-- $\mathfrak{m}_{1}(\mathfrak{m}(y,x),z)-\mathfrak{m}(\mathfrak{m}_{1}(y,x),z)+\mathfrak{m}_{1}(y,\mathfrak{m}(x,z))+\mathfrak{m}(y,\mathfrak{m}_{1}(x,z))=0.$
Or equivalently $\delta(\mathfrak{m}_{1})=[\mathfrak{m},\mathfrak{m}_{1}]_{MN}=0$.
The $1$-degree multiderivation $\mathfrak{m}_{1}$ is called the infinitesimal of the deformation $\mathfrak{m}_{t}$. More generally, if $\mathfrak{m}_{i}=0$ for $11\leq i\leq n-$ 1 and $\mathfrak{m}_{n}$ is non zero $1$-degree multiderivation then $\mathfrak{m}_{n}$ is called the $n$-infinitesimal of the deformation $\mathfrak{m}_{t}$. By the above discussion, the following proposition follows immediately.
Proposition 6.2.
The infinitesimal of the deformation $\mathfrak{m}_{t}$ is a 2-cocycle in $\mathfrak{D}^{1}(L)$. More generally, the $n$-infinitesimal is a 2-cocycle.
Now we give a notion of equivalence of two deformations. Let us denote a deformation $(L,A,\mathfrak{m}_{t},\ell_{t})$ of $(L,A,\mathfrak{m},\ell)$ simply by $L_{t}$. Let us consider two deformations $L_{t}$ and $L^{\prime}_{t}$ of $(L,A,\mathfrak{m},\ell)$, generated by $\mathfrak{m}_{i}$ and $\mathfrak{m}^{\prime}_{i}$, respectively, for $i\geq 0$.
Definition 6.3.
Two deformations $L_{t}$ and $L^{\prime}_{t}$ are said to be equivalent if there exists a formal automorphism
$\Phi_{t}:L[[t]]\rightarrow L[[t]]\text{ defined as }\Phi_{t}=id_{L}+\sum_{i\geq 1}t^{i}\phi_{i}$
where for each $i\geq 1$, $\phi_{i}:L\rightarrow L$ is a $\mathbb{K}$-linear map such that
[…]
Definition 6.4.
Any deformation that is equivalent to the deformation $\mathfrak{m}_{0}=\mathfrak{m}$ is said to be a trivial deformation.
Theorem 6.5.
The cohomology class of the infinitesimal of a deformation $\mathfrak{m}_{t}$ is determined by the equivalence class of $\mathfrak{m}_{t}$.
Proof 6.6.
Let $\Phi_{t}$ be an equivalence of deformation between $\mathfrak{m}_{t}$ and $\tilde{\mathfrak{m}}_{t}$. Then we get,
$\tilde{\mathfrak{m}}_{t}(x,y)=\Phi_{t}^{-1}\mathfrak{m}_{t}(\Phi_{t}x,\Phi_{t}y).$
Comparing the coefficients of $t$ from both sides of the above equation we have
$\tilde{\mathfrak{m}}_{1}(x,y)+\Phi_{1}(\mathfrak{m}_{0}(x,y))=\mathfrak{m}_{1}(x,y)+\mathfrak{m}_{0}(\Phi_{1}(x),y)+\mathfrak{m}_{0}(x,\Phi_{1}(y)),$
or equivalently,
$\mathfrak{m}_{1}-\tilde{\mathfrak{m}}_{1}=\delta(\phi_{1}).$
This establishes the result.
Definition 6.7.
A left-symmetric Rinehart algebra is said to be rigid if and only if every deformation of it is trivial.
Theorem 6.8.
A non-trivial deformation of a left-symmetric Rinehart algebra is equivalent to a deformation whose n-infinitesimal is not a coboundary for some $n\geq 1$.
Proof 6.9.
Let $\mathfrak{m}_{t}$ be a deformation of left-symmetric Rinehart algebra with n-infinitesimal $\mathfrak{m}_{n}$ for some $n\geq 1$. Assume that there exists a 2-cochain $\phi\in C^{1}(L,L)$ with $\delta(\phi)=\mathfrak{m}_{n}$. Then set
$\Phi_{t}=id_{L}+\phi t^{n}~{}~{}\mbox{and define}~{}~{}\bar{\mathfrak{m}_{t}}=\Phi_{t}\circ\mathfrak{m}_{t}\circ\Phi_{t}^{-1}.$
Then by computing the expression and comparing coefficients of $t^{n}$, we get
$\bar{\mathfrak{m}}_{n}-\mathfrak{m}_{n}=-\delta(\phi).$
So, ${\bar{\mathfrak{m}}_{n}}=0$. We can repeat the argument to kill off any infinitesimal, which is a coboundary.
Corollary 6.10.
If $H^{2}(L,L)=0$, then all deformations of $L$ are equivalent to a trivial deformation.
6.2 Obstructions to the extension theory of deformations
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra. Now we consider the problem of extending a deformation of $\mathfrak{m}$ of order $n$ to a deformation of $\mathfrak{m}$ of order $(n+1)$. Let $\mathfrak{m}_{t}$ and $\ell_{t}$ be a deformation of order $n$ of $\mathfrak{m}$ and $\ell$ respectively. That is
${\mathfrak{m}}_{t}=\sum^{n}_{i=0}\mathfrak{m}_{i}t^{i}=\mathfrak{m}+\sum^{n}_{i=1}\mathfrak{m}_{i}t^{i}\quad\text{and}\quad{\ell}_{t}=\sum^{n}_{i=0}\mathfrak{\ell}_{i}t^{i}=\ell+\sum^{n}_{i=1}\ell_{i}t^{i},$
where $\mathfrak{m}_{i}\in\mathfrak{D}^{1}(L)$ and $\ell_{i}:L\to Der(A)$ a linear map for each $1\leq i\leq n$ such that
$\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))=\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)-\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)),$ (20)
$\mathfrak{m}_{i}(ax,y)=a\mathfrak{m}_{i}(x,y),$ (21)
$\mathfrak{m}_{i}(x,ay)=a\mathfrak{m}_{i}(x,y)+\ell_{i}(x)ay$ (22)
for all $1\leq i,j\leq n$. If there exists a 2-cochain $\mathfrak{m}_{n+1}\in\mathfrak{D}^{1}(L)$ and $\ell_{n+1}:L\to Der(A)$ such that $(L,A,\tilde{\mathfrak{m}_{t}},\tilde{\ell_{t}})$ is a deformation of $(L,A,\mathfrak{m},\ell)$ of order $n+1$, where
$\tilde{\mathfrak{m}_{t}}=\mathfrak{m}_{t}+\mathfrak{m}_{n+1}t^{n+1}\quad\text{and}\quad\tilde{\ell_{t}}=\ell_{t}+\ell_{n+1}t^{n+1}.$
Then we say that $\mathfrak{m}_{t}$ extends to a deformation of order $(n+1)$. In this case $\mathfrak{m}_{t}$ is called extendable.
Definition 6.11.
Let $\mathfrak{m}_{t}$ be a deformation of $\mathfrak{m}$ of order $n$. Consider the cochain in $C^{3}(L,L)$ defined as
$\begin{split}Obs_{L}(x,y,z)=\sum_{\begin{subarray}{c}i+j=n+1;\\ i,j>0\end{subarray}}\Big{(}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}&(x,\mathfrak{m}_{j}(y,z))\\ &-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)))\Big{)},\end{split}$ (23)
for $x,y,z\in L$. The 3-cochain $Obs_{L}$ is called an obstruction cochain for extending the deformation of $\mathfrak{m}$ of order $n$ to a deformation of order $n+1$.
A straightforward computation gives the following
Proposition 6.12.
The obstructions are left-symmetric Rinehart algebra 3-cocycles.
Theorem 6.13.
Let $\mathfrak{m}_{t}$ be a deformation of $\mathfrak{m}$ of order $n$. Then $\mathfrak{m}_{t}$ extends to a deformation of order $n+1$ if and only if the cohomology class of $Obs_{L}$ vanishes.
Proof 6.14.
Suppose that a deformation $\mathfrak{m}_{t}$ of order $n$ extends to a deformation of order $n+1$. Then
$\sum_{\begin{subarray}{c}i+j=n+1;\\ i,j\geq 0\end{subarray}}\Big{(}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)))\Big{)}=0.$
As a result, we get $Obs_{L}=\delta(m_{n+1})$. So, the cohomology class of $Obs_{L}$ vanishes.
Conversely, let $Obs_{L}$ be a coboundary. Suppose that
$Obs_{L}=\delta(\mathfrak{m}_{n+1})$
for some 2-cochain $\mathfrak{m}_{n+1}$. Define a map $\tilde{\mathfrak{m}_{t}}:L[[t]]\times L[[t]]\rightarrow L[[t]]$ as follows
$\tilde{\mathfrak{m}_{t}}=\mathfrak{m}_{t}+\mathfrak{m}_{n+1}t^{n+1}.$
Then for any $x,y,z\in L$, the map $\tilde{\mathfrak{m}_{t}}$ satisfies the following identity
$\sum_{\begin{subarray}{c}i+j=n+1;\\ i,j\geq 0\end{subarray}}\Big{(}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)))\Big{)}=0.$
This, in turn, implies that $\tilde{\mathfrak{m}_{t}}$ is a deformation of $\mathfrak{m}$ extending $\mathfrak{m}_{t}$.
Corollary 6.15.
If $H^{3}(L,L)=0,$ then every 2-cocycle in $C^{2}(L,L)$ is the infinitesimal of some deformation of $\mathfrak{m}$.
6.3 Trivial deformation
We study deformations of left-symmetric Rinehart algebras using the deformation cohomology. Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra, and $\mathfrak{m}\in C^{2}(L,L)$. Consider a $t$-parameterized family of multiplications $\mathfrak{m}_{t}:L[[t]]\otimes L[[t]]\rightarrow L[[t]]$ and linear maps $\ell_{t}:L\to Der(A)$ given by
$\mathfrak{m}_{t}(x,y)=$ $x\cdot y+t\mathfrak{m}(x,y),$ (24)
$\ell_{t}=$ $\ell+t\Xi_{\mathfrak{m}}.$ (25)
If $L_{t}=(L,A,\mathfrak{m}_{t},\ell_{t})$ is a left-symmetric Rinehart algebra for all $t$, we say that $\mathfrak{m}$ generates a $1$-parameter infinitesimal deformation of $(L,A,\cdot,\ell)$
Since $\mathfrak{m}$ is a $2$-cochain, we have
$\mathfrak{m}(ax,y)=a\mathfrak{m}(x,y),\ \text{and}\ \mathfrak{m}(x,ay)=a\mathfrak{m}_{t}(x,y)+\Xi_{\mathfrak{m}}(x)(a)y,$
which implies that conditions (4) and (5) in Definition 3.1 are satisfied for $\mathfrak{m}_{t}$. Then we can deduce that $(L,A,\mathfrak{m}_{t},\ell_{t})$ is a deformation of $(L,A,\cdot,\ell)$ if and only if
$x\cdot\mathfrak{m}(y,z)-y\cdot\mathfrak{m}(x,z)+\mathfrak{m}(y,x)\cdot z-\mathfrak{m}(x,y)\cdot z$
$=\mathfrak{m}(y,x\cdot z)-\mathfrak{m}(x,y\cdot z)-\mathfrak{m}([x,y],z),$ (26)
and
$\mathfrak{m}(\mathfrak{m}(x,y),z)-\mathfrak{m}(x,\mathfrak{m}(y,z))$ $=$ $\mathfrak{m}(\mathfrak{m}(y,x),z)-\mathfrak{m}(y,\mathfrak{m}(x,z)).$ (27)
[…]
Recall that a deformation is said to be trivial if there exists a family of left-symmetric Rinehart algebra isomorphisms $\mathrm{Id}+tN:L_{t}\longrightarrow L$.
By direct computations, $L_{t}$ is trivial if and only if
$\mathfrak{m}(x,y)$ $=$ $x\cdot N(y)+N(x)\cdot y-N(x\cdot y),$ (28)
$N\mathfrak{m}(x,y)$ $=$ $N(x)\cdot N(y),$ (29)
$\ell\circ N$ $=$ $\Xi_{\mathfrak{m}}.$ (30)
[…]
$N(x)\cdot N(y)-x\cdot N(y)-N(x)\cdot y+N^{2}(x\cdot y)=0.$ (31)
Now we give the following definition.
Definition 6.16.
[…]
Obviously, any Nijenhuis operator on a left-symmetric Rinehart algebra is also a Nijenhuis operator on the corresponding sub-adjacent Lie-Rinehart algebra.
We have seen that a trivial deformation of a left-symmetric Rinehart algebra gives rise to a Nijenhuis operator. In fact, the converse is also true as can be seen from the following theorem.
Theorem 6.17.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ be a Nijenhuis operator. Then a deformation of $(L,A,\cdot,\ell)$ can be obtained by putting
$\mathfrak{m}(x,y)=\delta N(x,y).$
Furthermore, this deformation is trivial.
Proof 6.18.
Since $\mathfrak{m}$ is a coboundary, then it is a cocycle, i.e. equation (6.3) holds. To see that $\mathfrak{m}$ generates a deformation, we only need to show that (27) holds, which follows from the Nijenhuis condition (31). At the end, we can easily check that
$(\mathrm{Id}+tN)(x\cdot_{t}y)=(\mathrm{Id}+tN)(x)\cdot(\mathrm{Id}+tN)(y),\quad\ell\circ(\mathrm{Id}+tN)=\ell_{t},$
which implies that the deformation is trivial.
Theorem 6.19.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ be a Nijenhuis operator. Then $(L,A,\cdot_{N},\ell_{N}=\ell\circ N)$ is a left-symmetric Rinehart algebra, where
$x\cdot_{N}y=x\cdot N(y)+N(x)\cdot y-N(x\cdot y),\forall x,y\in L.$
Proof 6.20.
It is obvious to show that $(L,\cdot_{N})$ is a left-symmetric algebra and $\ell_{N}$ is a representation of L on $Der(A)$. Evidently, we have
$\ell_{N}(ax)=a\ell_{N}(x),\forall x\in L,a\in A.$
Furthermore, for any $x,y\in L$ and $a\in A$ we have
$x\cdot_{N}(ay)=$ $x\cdot N(ay)+N(x)\cdot(ay)-N(x\cdot(ay))$
$=$ $a(x\cdot N(y))+\ell(x)aN(y)+a(N(x)\cdot y)+\ell(N(x))ay-aN(x\cdot y)-N(\ell(x)ay)$
$=$ $a(x\cdot N(y)+N(x)\cdot y-N(x\cdot y))+\ell_{N}(x)ay+N(\ell(x)ay)-N(\ell(x)ay).$
$=$ $a(x\cdot_{N}y)+\ell_{N}(x)ay.$
Moreover,
$(ax)\cdot_{N}y=$ $(ax)\cdot N(y)+N(ax)\cdot y-N((ax)\cdot y)$
$=$ $a(x\cdot N(y))+a(N(x)\cdot y)-aN(x\cdot y)$
$=$ $a(x\cdot N(y)+N(x)\cdot y-N(x\cdot y))$
$=$ $a(x\cdot_{N}y).$
Then, $(L,A,\cdot_{N},\ell_{N}=\ell\circ N)$ is a a left-symmetric Rinehart algebra.
Immediately, we have the following result.
Lemma 6.21.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ be a Nijenhuis operator. Then for arbitrary positive $j,k\in\mathbb{N}$, the following equation holds
$N^{j}(x)\cdot N^{k}(y)-N^{k}(N^{j}(x)\cdot y)-N^{j}(x\cdot N^{k}(y))+N^{j+k}(x\cdot y)=0,\quad\forall~{}x,y\in L.$ (32)
If $N$ is invertible, this formula becomes valid for arbitrary $j,k\in\mathbb{Z}$.
By direct calculations, we have the following corollary.
Corollary 6.22.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ a Nijenhuis operator.
  • $\rm(i)$
    For all $k\in\mathbb{N}$, $(L,A,\cdot_{N^{k}},\ell_{N^{k}}=\ell\circ N^{k})$ is a left-symmetric Rinehart algebra.
  • $\rm(ii)$
    For all $l\in\mathbb{N}$, $N^{l}$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $(L,A,\cdot_{N^{k}},\ell_{N^{k}})$.
  • $\rm(iii)$
    The left-symmetric Rinehart algebras $(L,A,(\cdot_{N^{k}})_{N^{l}},\ell_{N^{k+l}})$ and $(L,A,\cdot_{N^{k+l}},\ell_{N^{k+l}})$ are the same.
  • $\rm(iv)$
    $N^{l}$ is a left-symmetric Rinehart algebra homomorphism from $(L,A,\cdot_{N^{k+l}},\ell_{N^{k+l}})$ to $(L,A,\cdot_{N^{k}},\ell_{N^{k}})$.
Theorem 6.23.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ be a Nijenhuis operator. Then the operator $P(N)=\sum_{i=0}^{n}c_{i}N^{i}$ is a Nijenhuis operator. If $N$ is invertible, then $Q(N)=\sum_{i=-m}^{n}c_{i}N^{i}$ is also a Nijenhuis operator.
Proof 6.24.
According to Lemma 6.21, we obtain, $\forall x,y\in L,$
$P(N)(x)\cdot P(N)(y)-P(N)(P(N)(x)\cdot y)-P(N)(x\cdot P(N)(y))+P^{2}(N)(x\cdot y)$
$=$ $\sum_{i,j=0}^{n}c_{j}c_{k}\Big{(}N^{j}(x)\cdot N^{k}(y)-N^{k}(N^{j}(x)\cdot y)-N^{j}(x\cdot N^{k}(y))+N^{j+k}(x\cdot y)\Big{)}=0,$
which implies that $P(N)$ is a Nijenhuis operator. Similarly we can easy check the second statement.
7 $\mathcal{O}$-operators and Nijenhuis operators
In this section, we highlight the relationships between $\mathcal{O}$-operators and Nijenhuis operators on left-symmetric Rinehart algebras. Moreover, we illustrate some connections between Nijenhuis operators and compatible $\mathcal{O}$-operators on left-symmetric Rinehart algebras.
7.1 Relationships between $\mathcal{O}$-operators and Nijenhuis operators
We first give the definitions of an $\mathcal{O}$-operator and of Rota-Baxter operator.
Definition 7.1.
An $\mathcal{O}$-operator on a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$ associated to a representation $(M;\rho,\mu)$ is a linear map $T:M\longrightarrow L$ satisfying
$T(au)$ $=$ $aT(u),$ (33)
$T(u)\cdot T(v)$ $=$ $T\Big{(}\rho(T(u))(v)+\mu(T(v))(u)\Big{)},\quad\forall u,v\in M,a\in A.$ (34)
Definition 7.2.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $\mathcal{R}:ker(\ell)\longrightarrow L$ a linear operator. If $\mathcal{R}$ satisfies
$\mathcal{R}(ax)$ $=$ $a\mathcal{R}(x),$ (35)
$\mathcal{R}(x)\cdot\mathcal{R}(y)$ $=$ $\mathcal{R}(\mathcal{R}(x)\cdot y+x\cdot\mathcal{R}(y)),\quad\forall x,y\in L,a\in A,$ (36)
then $\mathcal{R}$ is called a Rota-Baxter operator of weight 00 on $L$.
Notice that a Rota-Baxter operator of weight zero on a left-symmetric Rinehart algebra $L$ is exactly an $\mathcal{O}$-operator associated to the adjoint representation $(L;ad^{L},ad^{R})$.
The following proposition gives connections between Nijenhuis operators and Rota-Baxter operators.
Proposition 7.3.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N:L\longrightarrow L$ a linear operator.
  1. (i)
    If $N^{2}={\rm Id}$, then $N$ is a Nijenhuis operator if and only if $N\pm{\rm Id}$ is a Rota-Baxter operator of weight $\mp 2$ on $(L,A,\cdot,\ell)$.
  2. (ii)
    If $N^{2}=0$, then $N$ is a Nijenhuis operator if and only if $N$ is a Rota-Baxter operator of weight zero on $(L,A,\cdot,\ell)$.
  3. (iii)
    If $N^{2}=N$, then $N$ is a Nijenhuis operator if and only if $N$ is a Rota-Baxter operator of weight $-1$ on $(L,A,\cdot,\ell)$.
Proof 7.4.
For Item (i), for all $x,y\in L$ then we have
$(N-Id)$ $(x)\cdot(N-Id)(y)$
$\qquad-(N-Id)\big{(}(N-Id)(x)\cdot y+x\cdot(N-Id)(y)\big{)}+2(N-Id)(x\cdot y)$
$=N(x)\cdot N(y)-N\big{(}N(x)\cdot y+x\cdot N(y)\big{)}+x\cdot y$
$=N(x)\cdot N(y)-N\big{(}N(x)\cdot y+x\cdot N(y)\big{)}+N^{2}(x\cdot y).$
So $N$ is a Nijenhuis operator if and only if $N-Id$ is a Rota-Baxter operator of weight $2$ on $L$. Similarly, we obtain that $N$ is a Nijenhuis operator if and only if $N+Id$ is a Rota-Baxter operator of weight $-2$ on $L$.
Items (ii) and (iii) are obvious from the definitions of Nijenhuis operators and Rota-Baxter operator.
Proposition 7.5.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $(M;\rho,\mu)$ be a representation on $L$. Let $T:M\rightarrow L$ be a linear map. For any $\lambda$, $T$ is an $\mathcal{O}$-operator on $L$ associated to $(M;\rho,\mu)$ if and only if the linear map $\mathcal{R}_{T,\lambda}:=\left(\begin{array}[]{cc}0&T\\ 0&-\lambda{\rm Id}\end{array}\right)$ is a Rota-Baxter operator of weight $\lambda$ on the semidirect product left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$, where the multiplication $\cdot_{L\oplus M}$ is given by (9).
Proof 7.6.
It is easy to check the equation (35). Let $x_{1},x_{2}\in L$ and $m_{1},m_{2}\in M$,
$\mathcal{R}_{T,\lambda}(x_{1}+m_{1})\cdot_{L\oplus M}\mathcal{R}_{T,\lambda}(x_{2}+m_{2})$ $=(T(m_{1})-\lambda m_{1})\cdot_{L\oplus M}(T(m_{2})-\lambda m_{2})$
$=T(m_{1})\cdot T(m_{2})-\lambda\rho(T(m_{1})m_{2}-\lambda\mu(T(m_{2}))m_{1}.$ (37)
On the other hand,
$\mathcal{R}_{T,\lambda}\big{(}\mathcal{R}_{T,\lambda}(x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2})+(x_{1}+m_{1})\cdot_{L\oplus M}\mathcal{R}_{T,\lambda}(x_{2}+m_{2})\big{)}+$
$\quad\lambda\mathcal{R}_{T,\lambda}((x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2}))$
$=$ $\mathcal{R}_{T,\lambda}\big{(}(T(m_{1})-\lambda m_{1})\cdot_{L\oplus M}(x_{2}+m_{2})+(x_{1}+m_{1})\cdot_{L\oplus M}(T(m_{2})-\lambda m_{2})\big{)}+$
$\quad\lambda\mathcal{R}_{T,\lambda}(x_{1}\cdot x_{2}+\rho(x_{1})m_{2}+\mu(x_{2})m_{1})$
$=$ $\mathcal{R}_{T,\lambda}\big{(}T(m_{1})\cdot x_{2}+\rho(T(m_{1}))m_{2}-\lambda\mu(x_{2})m_{1}$ (38)
$\quad\qquad+x_{1}\cdot T(m_{2})-\lambda\rho(x_{1})m_{2}+\mu(T(m_{2}))m_{1}\big{)}$
$+\lambda\big{(}T(\rho(x_{1})m_{2})+T(\mu(x_{2})m_{1})-\lambda(\rho(x_{1})m_{2}+\mu(x_{2})m_{1})\big{)}$
$=$ $T\big{(}\rho(T(m_{1}))m_{2}+\mu(T(m_{2}))m_{1}\big{)}-\lambda T\big{(}\mu(x_{2})m_{1}+\rho(x_{1})m_{2}\big{)}$ (39)
$-\lambda\Big{(}\rho(T(m_{1}))m_{2}+\mu(T(m_{2}))m_{1}\Big{)}+\lambda^{2}\Big{(}\mu(x_{2})m_{1}+\rho(x_{1})m_{2}\big{)}$
$+\lambda\big{(}T(\rho(x_{1})m_{2})+T(\mu(x_{2})m_{1})\big{)}-\lambda^{2}\big{(}\rho(x_{1})m_{2}+\mu(x_{2})m_{1})\big{)}$
$=$ $T\big{(}\rho(T(m_{1}))m_{2}+\mu(T(m_{2}))m_{1}\big{)}-\lambda\rho(T(m_{1}))m_{2}-\lambda\mu(T(m_{2}))m_{1}.$ (40)
According to equations (37) and (40), $\mathcal{R}_{T,\lambda}$ is a Rota-Baxter operator of weight $\lambda$ on the semidirect product left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$ if and only if $T$ is an $\mathcal{O}$-operator on $(L,A,\cdot,\ell)$ associated to $(M;\rho,\mu)$.
Proposition 7.7.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and let $(M;\rho,\mu)$ be a representation on $L$. Let $T:M\rightarrow L$ be a linear map. Then the following statements are equivalent.
  1. (i)
    $T$ is an $\mathcal{O}$-operator on the left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$.
  2. (ii)
    $\mathcal{N}_{T}:=\left(\begin{array}[]{cc}0&T\\ 0&{\rm Id}\end{array}\right)$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$.
  3. (iii)
    $N_{T}:=\left(\begin{array}[]{cc}0&T\\ 0&0\end{array}\right)$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$.
Proof 7.8.
Note that $\mathcal{N}_{T}=\mathcal{R}_{T,-1}$ and $(\mathcal{N}_{T})^{2}=\mathcal{N}_{T}$, thus $\mathcal{N}_{T}$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$, using (iii) in Proposition 7.3.
Similarly $N_{T}=\mathcal{R}_{T,0}$ and $(N_{T})^{2}=0$, then $N_{T}$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$, according to (ii) in Proposition 7.3.
7.2 Compatible $\mathcal{O}$-operators and Nijenhuis operators
In this subsection we study compatibility of $\mathcal{O}$-operators and Nijenhuis operators. First we start with the following definition.
Definition 7.9.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and let $(M;\rho,\mu)$ be a representation. Let $T_{1},T_{2}:M\longrightarrow L$ be two $\mathcal{O}$-operators associated to $(M;\rho,\mu)$. Then $T_{1}$ and $T_{2}$ are called compatible if $T_{1}+T_{2}$ is an $\mathcal{O}$-operator associated to $(M;\rho,\mu)$.
Let $T_{1},T_{2}:M\longrightarrow L$ be two $\mathcal{O}$-operators on a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$ associated to a representation $(M;\rho,\mu)$ such that
$T_{1}(u)\cdot T_{2}(v)+T_{2}(u)\cdot T_{1}(v)$ $=T_{1}\Big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\Big{)}$
$\qquad+T_{2}\Big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\Big{)},$ (41)
for all $u,v\in M.$
Lemma 7.10.
Two operators $T_{1}$ and $T_{2}$ are compatible if and only if the equation (41) holds.
Proof 7.11.
For all $u,v\in M$, $a\in A$, we have
$(T_{1}+T_{2})(au)$ $=T_{1}(au)+T_{2}(au)$
$=aT_{1}(u)+aT_{2}(u)$
$=a(T_{1}+T_{2})(u).$
Furthermore,
$\!\!(T_{1}+T_{2})(u)\cdot(T_{1}+T_{2})(v)-(T_{1}+T_{2})\Big{(}\rho((T_{1}+T_{2})(u))(v)+\mu((T_{1}+T_{2})(v))(u)\Big{)}$
$=T_{1}(u)\cdot T_{1}(v)+T_{1}(u)\cdot T_{2}(v)+T_{2}(u)\cdot T_{1}(v)+T_{2}(u)\cdot T_{2}(v)$
$\qquad-(T_{1}+T_{2})\Big{(}\rho(T_{1}(u))(v)+\rho(T_{2}(u))(v)+\mu(T_{1}(v))(u)+\mu(T_{2}(v))(u)\Big{)}$
$=T_{1}(u)\cdot T_{1}(v)+T_{1}(u)\cdot T_{2}(v)+T_{2}(u)\cdot T_{1}(v)+T_{2}(u)\cdot T_{2}(v)$
$\qquad-T_{1}\big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\big{)}-T_{1}\big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\big{)}$
$\qquad-T_{2}\big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\big{)}-T_{2}\big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\big{)}$
$=T_{1}(u)\cdot T_{2}(v)+T_{2}(u)\cdot T_{1}(v)$
$\qquad-T_{1}\Big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\Big{)}-T_{2}\Big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\Big{)}$
Then $T_{1}+T_{2}$ is an $\mathcal{O}$-operator associated to $(M;\rho,\mu)$ if and only if equation (41) holds.
Remark 7.12.
Equation (41) implies that for any $k_{1},k_{2}$ the linear combination $k_{1}T_{1}+k_{2}T_{2}$ is an $\mathcal{O}$-operator.
There is a close relationship between a Nijenhuis operator and a pair of compatible $\mathcal{O}$-operators as can be seen from the following proposition.
Proposition 7.13.
Let $T_{1},T_{2}:M\longrightarrow L$ be two $\mathcal{O}$-operators on a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$ associated to a representation $(M;\rho,\mu)$. Suppose that $T_{2}$ is invertible. If $T_{1}$ and $T_{2}$ are compatible, then $N=T_{1}\circ T_{2}^{-1}$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$.
Proof 7.14.
For all $x,y\in L$, since $T_{2}$ is invertible, there exist $u,v\in M$ such that $T_{2}(u)=x,$ $T_{2}(v)=y$. Hence $N=T_{1}\circ T_{2}^{-1}$ is a Nijenhuis operator if and only if the following equation holds:
$NT_{2}(u)\cdot NT_{2}(v)=N(NT_{2}(u)\cdot T_{2}(v)+T_{2}(u)\cdot NT_{2}(v))-N^{2}(T_{2}(u)\cdot T_{2}(v)).$ (42)
Since $T_{1}=N\circ T_{2}$ is an $\mathcal{O}$-operator, the left-hand side of the above equation is
$NT_{2}(\rho(NT_{2}(u))(v)+\mu(NT_{2}(v))(u)).$
Using the fact that $T_{2}$ and $T_{1}=N\circ T_{2}$ are two compatible $\mathcal{O}$-operators, we get
$NT_{2}(u)\cdot T_{2}(v)+T_{2}(u)\cdot NT_{2}(v)$
$=$ $T_{2}(\rho(NT_{2}(u))(v)+\mu(NT_{2}(v))(u))+NT_{2}(\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u))$
$=$ $T_{2}(\rho(NT_{2}(u))(v)+\mu(NT_{2}(v))(u))+N(T_{2}(u)\cdot T_{2}(v)).$
Hence, equation (42) holds by acting $N$ on both sides of the last equality.
Using an $\mathcal{O}$-operator and a Nijenhuis operator, we can construct a pair of compatible $\mathcal{O}$-operators.
Proposition 7.15.
Let $T:M\longrightarrow L$ be an $\mathcal{O}$-operator on a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$ associated to a representation $(M;\rho,\mu)$ and let $N$ be a Nijenhuis operator on $(L,A,\cdot,\ell)$. Then $N\circ T$ is an $\mathcal{O}$-operator on the left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$ associated to $(M;\rho,\mu)$ if and only if for all $u,v\in M$, the following equation holds:
$N\Big{(}NT(u)\cdot T(v)+T(u)\cdot NT(v)\Big{)}$
$=$ $N\Big{(}T\big{(}\rho(NT(u))(v)+\mu(NT(v))(u)\big{)}+NT\big{(}\rho(T(u))(v)+\mu(T(v))(u)\big{)}\Big{)}.$ (43)
If in addition $N$ is invertible, then $T$ and $NT$ are compatible. More explicitly, for any $\mathcal{O}$-operator $T$, if there exists an invertible Nijenhuis operator $N$ such that $NT$ is also an $\mathcal{O}$-operator, then $T$ and $NT$ are compatible.
Proof 7.16.
Let $u,v\in M$ and $a\in A$, we have
$NT(au)=N(T(au))=N(aT(u))=aNT(u).$
In addition, since $N$ is a Nijenhuis operator and $T$ is an $\mathcal{O}$-operator we have
$NT(u)\cdot NT(v)$ $=N\Big{(}NT(u)\cdot T(v)+T(u)\cdot NT(v)\Big{)}-N^{2}(T(u)\cdot T(v))$
$=NT\Big{(}\rho(NT(u))(v)+\mu(NT(v))(u)\Big{)}$
if and only if (43) holds.
If $NT$ is an $\mathcal{O}$-operator and $N$ is invertible, then we have
$T(u)\cdot T(v)+T(u)\cdot NT(v)$
$=T\big{(}\rho(NT(u))(v)+\mu(NT(v))(u)\big{)}+NT\big{(}\rho(T(u))(v)+\mu(T(v))(u)\big{)},$
which is exactly the condition that $NT$ and $T$ are compatible.
The following result is an immediate consequence of the last two propositions.
Corollary 7.17.
Let $T_{1},T_{2}:M\longrightarrow L$ be two $\mathcal{O}$-operators on a left-symmetric Rinheart algebra $(L,A,\cdot,\ell)$ associated to a representation $(M;\rho,\mu)$. Suppose that $T_{1}$ and $T_{2}$ are invertible. Then $T_{1}$ and $T_{2}$ are compatible if and only if $N=T_{1}\circ T_{2}^{-1}$ is a Nijenhuis operator.
Acknowledgement
The authors are thankful to the Deanship of Graduate Studies and Scientific Research at University of Bisha for supporting this work through the Fast-Track Research Support Program. Mohamed Elhamdadi was partially supported by Simons Foundation collaboration grant 712462.